Solve this system by the addition method:
x/4 + y/3 =5
x - y =6

Answer by stanbon(75887)

Solve this system by the addition method:
x/4 + y/3 =5
x - y =6
--------------------
Multiply thru the 1st equation by 12
Multiply thru the 2nd equation by -3
-----------------------------------------
3x + 4y = 60
-3x+ 3y =-18
-------------------
Add and solve for "y":
7y = 42
y = 6
------------------------------
Substitute into x-y=6 to solve for "x":
x - 6 = 6
x = 12
---------------------------------------
Cheers,
Stan H.